# Tallahassee, FL 32306-4350, USA

### v1

@@ 4.1 Results for the $U(3)^{5}$ flavour symmetric SMEFT
We now discuss the results of the global fit for the $U(3)^{5}$ assumption, whose results are presented in Table LABEL:tab:u3-glob and Figure 3. It is well known that, in the Warsaw basis, there are flat directions in the fit to EW precision observables in terms of SMEFT coefficients at the EW scale, see e.g. ref. Falkowski:2014tna . These flat directions are lifted by the inclusion of other observables, for example in the Higgs sector, but leave their footprint in terms of correlations among the coefficients. Furthermore, the flat directions at the EW scale get distorted by the RGE. While the full correlation matrix is too large to be presented here, we can mention a few examples of highly correlated coefficients. We observe a correlation of $0.97$ between $C_{\phi B}$ and $C_{\phi WB}$, of $0.90$ between $C_{\phi WB}$ and $C_{\phi D}$, of $-0.88$ between $C_{\phi B}$ and $C_{\phi D}$, and of $-0.83$ between $C_{\phi W}$ and $C_{\phi WB}$. Those correlations are induced by EW and Higgs observables. Moreover, we observe a strong correlation of $-0.998$ between $C_{ll}^{[aabb]}$ and $C_{ee}$, of $0.84$ between $C_{ll}^{[aabb]}$ and $C_{ll}^{[abba]}$, and of $-0.87$ between $C_{ll}^{[abba]}$ and $C_{ee}$, all induced by EW observables, in particular, two-to-two fermion processes from LEP2. We also observe a correlation of $0.88$ between $C_{\phi G}$ and $C_{G}$, due to Higgs observables, and of $-0.83$ between $C_{qu}^{(1)}$ and $C_{qu}^{(8)}$, which is reasonably aligned with the expected RGE contributions from these two operators to ${\cal O}_{u\phi}^{[33]}$, whose coefficient modifies the top-quark Yukawa coupling. Finally, we mention a correlation of $-0.80$ between $C_{lq}^{(1)}$ and $C_{eu}$, due to DY observables. In spite of the dilution of the constraints due to interference between the coefficients, as can be seen from Figure 3 and Table LABEL:tab:u3-glob, we still obtain $\Lambda/\sqrt{|C|}\gtrsim 11$ TeV for $C_{\phi G}$, $\Lambda/\sqrt{|C|}\gtrsim 5$ TeV for $C_{lq}^{(3)}$, and $\Lambda/\sqrt{|C|}\gtrsim 4$ TeV for $C_{\phi q}^{(3)}$, while several other coefficients give bounds of the order of $2$-$3$ TeV. Therefore, for a generic strongly-interacting $U(3)^{5}$ flavour symmetric theory, current data constrain NP to be well above the TeV scale, making the SMEFT approach fully consistent, while weakly-interacting $U(3)^{5}$ flavour symmetric theories are still allowed at the TeV scale, where however the SMEFT approach might not be fully consistent for high-$p_{T}$ observables, and direct searches could be more effective.
@@
Finally, regarding the relaxation of the bounds in going from the individual fits to the global one due to mutual strong correlations, we would like to emphasise that this implies that a lower effective NP scale $\Lambda/\sqrt{|C|}$ with respect to the one indicated by individual fits is viable only for those UV completions that incorporate the phenomenologically required correlations. Conversely, we note that taking into account the theory prior imposing perturbativity on the Wilson coefficients actually tightens many bounds. Table LABEL:tab:u3-glob shows that several limits on $\Lambda/\sqrt{\lvert C\rvert}$ at 10 TeV now exceed those at 3 TeV. In particular, although correlations with other operators substantially weaken the global-fit constraints on $C_{\phi B}$ and $C_{lq}^{(3)}$, restricting the allowed ranges of the correlated $C/\Lambda^{2}$ coefficients ensures that the bounds on $\Lambda/\sqrt{\lvert C_{\phi B}\rvert}$ and $\Lambda/\sqrt{\lvert C_{lq}^{(3)}\rvert}$ remain comparably strong at 10 TeV.
@@ 4.2 Results for the $U(2)^{5}$ flavour symmetric SMEFT
The results of the individual fits for the $U(2)^{5}$ flavour symmetric SMEFT in the UP basis are reported in Table LABEL:tab:u2-ind and in Figures 4-7, which have the same structure as Table LABEL:tab:u3-ind and Figure 1 discussed in Section 4.1 for the $U(3)^{5}$ case, to which we refer. The limits on $\Lambda/\sqrt{|C|}$ are also summarised in Figure 8, where we only show the coefficients that can be constrained individually within the pertubative regime in the case of $\Lambda=3$ TeV. Furthermore, a comparison between the case of the UP and DOWN bases is summarised in Table LABEL:tab:u2-up-down and Figure 9 for the case of those coefficients that are affected by the choice of flavour basis. Finally, the global fit results are shown in Table LABEL:tab:u2-glob and Figures 10 and 11 where we show the results for the case of $\Lambda=10$ TeV.
@@
Furthermore, as shown in Table LABEL:tab:u2-up-down and Figure 9, the choice of flavour basis has a pronounced impact on the mentioned $b$-quark dipole operators. Aligning to the DOWN basis, as expected, strongly suppresses contributions to $b$-quark flavour transitions, relaxing the bound on ${C}_{dB}^{[33]}$ by nearly two orders of magnitude. Even in this case, however, flavour observables continue to provide the most stringent constraints on these operators. This pattern extends, to varying degrees, to most of the operators shown in Figure 9: bounds in the DOWN basis are systematically weaker due to flavour-space alignment and the precision of $B$-physics measurements. A few notable exceptions are the coefficients ${C}_{\phi q}^{(1)}$ and ${C}_{\phi q}^{(3)}$, for which flavour-diagonal electroweak observables impose similarly strong limits.
@@
We now discuss the results of the global fits for the $U(2)^{5}$ assumption, choosing as a representative example the results obtained with $\Lambda=10$ TeV in the UP basis shown in Table LABEL:tab:u2-glob and Figures 10 and 11. We begin by highlighting the most important correlations between the different Wilson coefficients. Let us start with the coefficients mainly constrained by Higgs-boson and EW observables. For example, the coefficient $C_{\phi G}$ becomes degenerate with $C_{uG}^{[33]}$, with a correlation coefficient of $-0.94$. However, imposing the perturbativity prior on $C_{uG}^{[33]}$, one still gets a strong constraint on $C_{\phi G}$, with a $95\%$ HPDI interval of $[-0.32,1.62]$ corresponding to a bound on $\Lambda/\sqrt{|C_{\phi G}|}$ of $8$ TeV, to be compared with the bound of $19$ TeV obtained in the individual fit (see Table LABEL:tab:u2-ind). A similar phenomenon occurs for other pairs of coefficients, such as $C_{e\phi}^{[33]}$ and $C_{lequ}^{(1)[3333]}$, with a correlation coefficient of $0.87$ and $C_{lequ}^{(1)[3333]}$ hitting the perturbativity limit; $C_{d\phi}^{[33]}$ and $C_{quqd}^{(1)[3333]}$, with a correlation coefficient of $-0.83$ and $C_{quqd}^{(1)[3333]}$ hitting the perturbativity limit; $C_{dB}^{[33]}$ and $C_{dW}^{[33]}$, with a correlation coefficient of $0.99$, and $C_{dW}^{[33]}$ hitting the perturbativity limit; $C_{ll}^{[aabb]}$ and $C_{ee}^{[aabb]}$, with a correlation coefficient of $-0.9$ and both coefficients hitting the perturbativity limit. Concerning four fermion operators, one faces more complex interference patterns involving several coefficients. For example, the coefficient $C_{qq}^{(3)[aabb]}$ is strongly correlated with $C_{lq}^{(3)[aabb]}$ and $C_{lq}^{(3)[33aa]}$, with correlation coefficients of $0.86$ and $0.81$, respectively. Therefore, a large set of four fermions operators cannot be constrained in the global fit with current data.
@@
A direct comparison between the global fit results in the $U(2)^{5}$ and $U(3)^{5}$ flavour-symmetric scenarios can be made by examining the last column of Tables LABEL:tab:u2-glob and LABEL:tab:u3-glob. We have already discussed how certain dipole interactions involving the bottom quark (${C}_{dB,dW,dG}^{[33]}$), which are absent under the $U(3)^{5}$ assumption, receive strong constraints from flavour observables in the $U(2)^{5}$ framework. On the other hand, when comparing the global bounds on operators such as ${C}_{\phi q}^{(1)}$ and ${C}_{\phi q}^{(3)}$, which modify the neutral and charged current couplings of left-handed quarks, we find that the $U(2)^{5}$ symmetry still offers significant protection against flavour constraints, yielding bounds comparable to those in the $U(3)^{5}$ scenario. However, this does not hold for individual fits. In particular, interactions involving the left-handed third-generation quarks are more strongly constrained by flavour observables, while those involving the light generations remain primarily constrained by electroweak data.  [^16] Interestingly, and as also pointed out in other studies, for the right-handed interactions with the top-quark, ${C}_{\phi u}^{[33]}$, the strongest individual bound comes from electroweak precision observables as a result of mixing between this operator and the custodial symmetry breaking operator ${C}_{\phi D}$ deBlas:2015aea (see also Stefanek:2024kds ). A similar effect is induced by the left-handed counterpart ${C}_{\phi q}^{{(1})[33]}$, making electroweak bounds comparable to the flavour ones in this case (see Table LABEL:tab:u2-ind).
Similar conclusions can be derived by looking at lepton-quark left-handed operators, where the separation of the third and first two families in the $U(2)^{5}$ also helps to identify the leading source of the bound in the $U(3)^{5}$-symmetric case. (For both the ${C}_{lq}^{(1),(3)}$ operators, this comes from the entries involving only the first two quark families, where Drell-Yan at high energies still seems to be the leading constraint.) Unfortunately, the lack of constraints for many of the interactions in the $U(2)^{5}$ global fit does not allow to derive further information from the comparison with the $U(3)^{5}$ case.
@@ 5 Conclusions
In the $U(3)^{5}$-symmetric case, the constraints on individual Wilson coefficients translate into lower bounds on the effective NP scale $\Lambda/\sqrt{|C_{i}|}$ reaching values up to $\sim 25$ TeV, particularly when RGE effects are included. We observe that the running of SMEFT operators plays a crucial role in tightening constraints for operators such as $\mathcal{O}_{G}$ and $\mathcal{O}_{W}$, which gain sensitivity through mixing into observables in the Higgs and EW sectors. In global fits – where all relevant Wilson coefficients are varied simultaneously under the assumed flavour symmetry – strong correlations among them tend to relax many of the individual bounds. For instance, large correlations arise among several bosonic and four-fermion operators. Nevertheless, meaningful bounds persist: $\Lambda/\sqrt{|C_{\phi G}|}\gtrsim 11$ TeV, $\Lambda/\sqrt{|C_{lq}^{(3)}|}\gtrsim 5$ TeV, and $\Lambda/\sqrt{|C_{\phi q}^{(3)}|}\gtrsim 4$ TeV, while other coefficients yield bounds in the 22223333 TeV range. These findings indicate that generic strongly-coupled $U(3)^{5}$-symmetric NP scenarios are pushed above the TeV scale, ensuring the internal consistency of the SMEFT expansion. Conversely, weakly-coupled UV completions may still reside near the TeV scale, though in such cases the SMEFT description must be applied with care, particularly for high-$p_{T}$ observables where EFT validity can break down.
In the $U(2)^{5}$ case, the orientation in flavour space becomes relevant, and the inclusion of flavour observables has a decisive impact on the analysis. In this study, we considered two complementary cases: one in which NP effects are introduced in the basis where the up-quark Yukawa matrix is diagonal – the UP basis – and another where NP effects are defined in the basis with diagonal down-quark Yukawas, the DOWN basis. In both cases, flavour misalignment stems from the left-handed quark sector, while the right-handed quark Yukawa couplings and the charged-lepton Yukawa matrix are flavour diagonal.
@@ Acknowledgements
V.M. is grateful to Marcel Vos for insightful discussions during their joint work. He also thanks Antonio Jaques Costa for valuable advice on several experimental measurements. The work of J.B. has been partially funded by the FEDER/Junta de Andalucía project grant P18-FRJ-3735, MICIU/AEI/10.13039/501100011033 and FEDER/UE (grant PID2022-139466NB-C21). The work of V.M. is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 949451) and a Royal Society University Research Fellowship through grant URF/R1/201553. The work of L.R. is supported in part by the U.S. Department of Energy under grant DE- SC0010102 and by INFN through a Foreign Visiting Scientist Fellowship. This work was supported in part by the European Union - Next Generation EU under italian MUR grant PRIN-2022-RXEZCJ. L.S. and M.V. acknowledge support from the project “Exploring New Physics” funded by INFN.

### v2

@@ 4.1 Results for the $U(3)^{5}$ flavour symmetric SMEFT
We now discuss the results of the global fit for the $U(3)^{5}$ assumption, whose results are presented in Table LABEL:tab:u3-glob and Figure 3. It is well known that, in the Warsaw basis, there are flat directions in the fit to EW precision observables in terms of SMEFT coefficients at the EW scale, see e.g. ref. Falkowski and Riva (2015). These flat directions are lifted by the inclusion of other observables, for example in the Higgs sector, but leave their footprint in terms of correlations among the coefficients. Furthermore, the flat directions at the EW scale get distorted by the RGE. While the full correlation matrix is too large to be presented here, we can mention a few examples of highly correlated coefficients. We observe a correlation of $0.95$ between $C_{\phi B}$ and $C_{\phi WB}$, of 0.92-0.92 between $C_{\phi WB}$ and $C_{\phi D}$, of 0.89-0.89 between $C_{\phi B}$ and $C_{\phi D}$, of 0.85-0.85 between $C_{\phi D}$ and $C_{\phi e}$, of $0.84$ between $C_{\phi l}^{(1)}$ and $C_{\phi e}$, and of $0.83$ between $C_{\phi W}$ and $C_{\phi WB}$. Those correlations are induced by EW and Higgs observables. Moreover, we observe a strong correlation of 0.998-0.998 between $C_{ll}^{[aabb]}$ and $C_{ee}$, of $0.83$ between $C_{ll}^{[aabb]}$ and $C_{ll}^{[abba]}$, and of 0.86-0.86 between $C_{ll}^{[abba]}$ and $C_{ee}$, all induced by EW observables, in particular, two-to-two fermion processes from LEP2. We also observe a correlation of $0.88$ between $C_{\phi G}$ and $C_{G}$, due to Higgs observables, and of 0.83-0.83 between $C_{qu}^{(1)}$ and $C_{qu}^{(8)}$, which is reasonably aligned with the expected RGE contributions from these two operators to ${\cal O}_{u\phi}^{[33]}$, whose coefficient modifies the top-quark Yukawa coupling. Finally, we mention a correlation of 0.87-0.87 between $C_{lq}^{(1)}$ and $C_{eu}$, due to DY observables. In spite of the dilution of the constraints due to interference between the coefficients, as can be seen from Figure 3 and Table LABEL:tab:u3-glob, we still obtain $\Lambda/\sqrt{|C_{i}|}\gtrsim 14$ TeV for $C_{\phi G}$, $\Lambda/\sqrt{|C_{i}|}\gtrsim 5$ TeV for $C_{lq}^{(3)}$, while several other coefficients give bounds of the order of 22-33 TeV. Therefore, for a generic strongly-interacting $U(3)^{5}$ flavour symmetric theory, current data constrain NP to be well above the TeV scale, making the SMEFT approach fully consistent, while weakly-interacting $U(3)^{5}$ flavour symmetric theories are still allowed at the TeV scale, where however the SMEFT approach might not be fully consistent for high-$p_{T}$ observables, and direct searches could be more effective.
@@
Finally, regarding the relaxation of the bounds in going from the individual fits to the global one due to mutual strong correlations, we would like to emphasise that this implies that a lower effective NP scale $\Lambda/\sqrt{|C_{i}|}$ with respect to the one indicated by individual fits is viable only for those UV completions that incorporate the phenomenologically required correlations. Conversely, we note that taking into account the theory prior imposing perturbativity on the Wilson coefficients actually tightens many bounds. Table LABEL:tab:u3-glob shows that several limits on $\Lambda/\sqrt{\lvert C_{i}\rvert}$ at 10 TeV now exceed those at 3 TeV. In particular, although correlations with other operators substantially weaken the global-fit constraints on $C_{\phi B}$ and $C_{lq}^{(3)}$, restricting the allowed ranges of the correlated $C_{i}/\Lambda^{2}$ coefficients ensures that the bounds on $\Lambda/\sqrt{\lvert C_{\phi B}\rvert}$ and $\Lambda/\sqrt{\lvert C_{lq}^{(3)}\rvert}$ remain comparably strong at 10 TeV.
@@ 4.2 Results for the $U(2)^{5}$ flavour symmetric SMEFT
The results of the individual fits for the $U(2)^{5}$ flavour symmetric SMEFT in the UP basis are reported in Table LABEL:tab:u2-ind and in Figures 4-7, which have the same structure as Table LABEL:tab:u3-ind and Figure 1 discussed in Section 4.1 for the $U(3)^{5}$ case, to which we refer. The limits on $\Lambda/\sqrt{|C_{i}|}$ are also summarised in Figure 8, where we only show the coefficients that can be constrained individually within the pertubative regime in the case of $\Lambda=3$ TeV. Furthermore, a comparison between the case of the UP and DOWN bases is summarised in Table LABEL:tab:u2-up-down and Figure 9 for the case of those coefficients that are affected by the choice of flavour basis. Finally, the global fit results are shown in Table LABEL:tab:u2-glob and Figures 10 and 11 where we show the results for the case of $\Lambda=10$ TeV.
@@
Furthermore, as shown in Table LABEL:tab:u2-up-down and Figure 9, the choice of flavour basis has a pronounced impact on the mentioned $b$-quark dipole operators. Aligning to the DOWN basis, as expected, strongly suppresses contributions to $b$-quark flavour transitions, relaxing the bound on ${C}_{dB}^{[33]}$ by nearly two orders of magnitude. Even in this case, however, flavour observables continue to provide the most stringent constraints on these operators. This pattern extends, to varying degrees, to most of the operators shown in Figure 9: bounds in the DOWN basis are systematically weaker due to flavour-space alignment and the precision of $B$-physics measurements.
@@
We now discuss the results of the global fits for the $U(2)^{5}$ assumption, choosing as a representative example the results obtained with $\Lambda=10$ TeV in the UP basis shown in Table LABEL:tab:u2-glob and Figures 10 and 11. We begin by highlighting the most important correlations between the different Wilson coefficients. Let us start with the coefficients mainly constrained by Higgs-boson and EW observables. For example, the coefficient $C_{\phi G}$ becomes degenerate with $C_{uG}^{[33]}$, with a correlation coefficient of 0.91-0.91. However, imposing the perturbativity prior on $C_{uG}^{[33]}$, one still gets a strong constraint on $C_{\phi G}$, with a $95\%$ HPDI interval of [0.30,0.85][-0.30,0.85] corresponding to a bound on $\Lambda/\sqrt{|C_{\phi G}|}$ of $10.5$ TeV, to be compared with the bound of $19$ TeV obtained in the individual fit (see Table LABEL:tab:u2-ind). A similar phenomenon occurs for other pairs of coefficients, such as $C_{e\phi}^{[33]}$ and $C_{lequ}^{(1)[3333]}$, with a correlation coefficient of $0.91$ and $C_{lequ}^{(1)[3333]}$ hitting the perturbativity limit; $C_{d\phi}^{[33]}$ and $C_{quqd}^{(1)[3333]}$, with a correlation coefficient of 0.91-0.91 and $C_{quqd}^{(1)[3333]}$ hitting the perturbativity limit; $C_{dB}^{[33]}$ and $C_{dW}^{[33]}$, with a correlation coefficient of $0.99$, and $C_{dW}^{[33]}$ hitting the perturbativity limit; $C_{ll}^{[aabb]}$ and $C_{ee}^{[aabb]}$, with a correlation coefficient of 0.82-0.82 and both coefficients hitting the perturbativity limit. Another pair of strongly correlated coefficients is given by $C_{\phi e}^{[33]}$ and $C_{\phi e}^{[aa]}$, with a correlation coefficient of $0.84$. Concerning four fermion operators, one faces more complex interference patterns involving several coefficients. For example, the coefficient $C_{qq}^{(3)[aabb]}$ is strongly correlated with $C_{lq}^{(3)[aabb]}$ which is in turn strongly correlated with $C_{lq}^{(3)[33aa]}$, with correlation coefficients of $0.93$ and $0.85$, respectively. Therefore, a large set of four fermions operators cannot be constrained in the global fit with current data.
@@
A direct comparison between the global fit results in the $U(2)^{5}$ and $U(3)^{5}$ flavour-symmetric scenarios can be made by examining the last column of Tables LABEL:tab:u2-glob and LABEL:tab:u3-glob. We have already discussed how certain dipole interactions involving the bottom quark (${C}_{dB,dW,dG}^{[33]}$), which are absent under the $U(3)^{5}$ assumption, receive strong constraints from flavour observables in the $U(2)^{5}$ framework. On the other hand, when comparing the global bounds on operators such as ${C}_{\phi q}^{(1)}$ and ${C}_{\phi q}^{(3)}$, which modify the neutral and charged current couplings of left-handed quarks, we find that the $U(2)^{5}$ symmetry still offers significant protection against flavour constraints, yielding bounds comparable to those in the $U(3)^{5}$ scenario. However, this does not hold for individual fits. In particular, interactions involving the left-handed third-generation quarks are more strongly constrained by flavour observables, while those involving the light generations remain primarily constrained by electroweak data. [^17] Interestingly, and as also pointed out in other studies, for the right-handed interactions with the top-quark, ${C}_{\phi u}^{[33]}$, the strongest individual bound comes from electroweak precision observables as a result of mixing between this operator and the custodial symmetry breaking operator ${C}_{\phi D}$ de Blas et al. (2015) (see also Stefanek (2024)). A similar effect is induced by the left-handed counterpart ${C}_{\phi q}^{{(1})[33]}$, making electroweak bounds comparable to the flavour ones in this case (see Table LABEL:tab:u2-ind).
Similar conclusions can be derived by looking at lepton-quark left-handed operators, where the separation of the third and first two families in the $U(2)^{5}$ also helps to identify the leading source of the bound in the $U(3)^{5}$-symmetric case. For both the ${C}_{lq}^{(1),(3)}$ operators, this comes from the entries involving only the first two quark families, where Drell-Yan at high energies still seems to be the leading constraint. Unfortunately, the lack of constraints for many of the interactions in the $U(2)^{5}$ global fit does not allow to derive further information from the comparison with the $U(3)^{5}$ case.
@@ 5 Conclusions
In the $U(3)^{5}$-symmetric case, the constraints on individual Wilson coefficients translate into lower bounds on the effective NP scale $\Lambda/\sqrt{|C_{i}|}$ reaching values up to $\sim 25$ TeV, particularly when RGE effects are included. We observe that the running of SMEFT operators plays a crucial role in tightening constraints for operators such as $\mathcal{O}_{G}$ and $\mathcal{O}_{W}$, which gain sensitivity through mixing into observables in the Higgs and EW sectors. In global fits – where all relevant Wilson coefficients are varied simultaneously under the assumed flavour symmetry – strong correlations among them tend to relax many of the individual bounds. For instance, large correlations arise among several bosonic and four-fermion operators. Nevertheless, meaningful bounds persist: $\Lambda/\sqrt{|C_{\phi G}|}\gtrsim 14$ TeV and $\Lambda/\sqrt{|C_{lq}^{(3)}|}\gtrsim 5$ TeV, while other coefficients yield bounds in the 2233 TeV range. These findings indicate that generic strongly-coupled $U(3)^{5}$-symmetric NP scenarios are pushed above the TeV scale, ensuring the internal consistency of the SMEFT expansion. Conversely, weakly-coupled UV completions may still reside near the TeV scale, though in such cases the SMEFT description must be applied with care, particularly for high-$p_{T}$ observables where EFT validity can break down.
In the $U(2)^{5}$ case, the orientation in flavour space becomes relevant, and the inclusion of flavour observables has a decisive impact on the analysis. In this study, we considered two complementary cases: one in which NP effects are introduced in the basis where the up-quark Yukawa matrix is diagonal – the UP basis – and another where NP effects are defined in the basis with diagonal down-quark Yukawas, the DOWN basis. In both cases, flavour misalignment stems from the left-handed quark sector, while the right-handed quarks can always be aligned with the direction of flavour-diagonal Yukawa couplings. In the lepton sector, the charged-lepton flavours can also always be rotated such that their Yukawa interactions are flavour diagonal.
@@ Acknowledgements
V.M. is grateful to Marcel Vos for insightful discussions during their joint work. He also thanks Antonio Jaques Costa for valuable advice on several experimental measurements. We thank P. Stangl for useful comments on the manuscript. The work of J.B. has been partially funded by the FEDER/Junta de Andalucía project grant P18-FRJ-3735, MICIU/AEI/10.13039/501100011033 and FEDER/UE (grant PID2022-139466NB-C21). The work of V.M. is supported by the MICIU through a Beatriz Galindo Junior grant (BG24/00038), the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 949451) and a Royal Society University Research Fellowship through grant URF/R1/201553. The work of L.R. is supported in part by the U.S. Department of Energy under grant DE-SC0010102 and by INFN through a Foreign Visiting Scientist Fellowship. This work was supported in part by the European Union - Next Generation EU under italian MUR grant PRIN-2022-RXEZCJ. L.S. and M.V. acknowledge support from the project “Exploring New Physics” funded by INFN.